OpenStudy (anonymous):

Enter the quadratic equation in standard form. x+10=3(x-1)^2

OpenStudy (anonymous):

you foil the (x-1) and multiply that answer by 3 to get x+10=3x^2-6x=3. then bring over the x and the 10 to get 3x^2-7x-7=0

OpenStudy (anonymous):

Thank you!